# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### 29 Nov 2023 v1

@@ 10 More Experimental Results
Figure 7: Qualitative results on HuMMan [3]. ”Input” refers to input images of the model. ”Overlay” means we project the recovered mesh back onto the original images. ”Cam” and ”World” show the output in camera and world coordinates, respectively.
Design Details To find the optimal design for our model, we conduct further ablation experiments. Firstly, we test different downsampling methods. Evidently, using markers for downsampling leads to better learning of human body features. Here, ”markers” refer to mocap markers. The improved performance might be because mocap markers serve as a better proxy for inferring human body information [55]. However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.
@@
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of ”2 param$+$1 vert” has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
Qualitative Results More visualizations are generated to give a more comprehensive reference. Also, we welcome running our codes on your samples. Back to qualitative results, from Fig. 9, we can see that our model is able to recover the human body very well when facing various scenes with various characters in various poses. We purposely show the results from multiple angles and in world space. In the past, many methods can also make the reconstructed mesh match the original image perfectly. But the reconstructed body doesn’t look good if we change the visual angle. Moreover, if we put it in the world space, the pose would behave abnormally.
It may not be noticeable from Fig. 9 because they fit the weak-perspective camera model and the assumption of $R_{c}=I$. But from Fig. 7, the drawbacks of the traditional method are undeniable. The reconstructed human body is entirely lopsided in world coordinate. But with our OrientCorrect module, the pose can be corrected very well to achieve a reasonable pose recovery in world coordinate. This method has a wide range of applications, allowing the model to reconstruct a proper human body when facing in-the-wild images. Of course, W-HMR is still not a perfect model, and for this reason, we show the failure cases in Fig. 8. We can find that our model is still not able to achieve accurate reconstruction when facing extreme poses (yoga, parkour), which may be due to the lack of relevant training data or may be limited by the iterative method. Secondly, when the human limbs are close to the body, there is often through-molding, which requires an extra optimization-based method to improve it or a human model with a higher degree of freedom. These will be one of the directions worth researching in the future.
Figure 8: Failure cases on unusual samples. Although our model performs well in most cases, W-HMR still fails in some extreme situations. Such as, the human makes difficult poses, the body parts are hard to distinguish from the background, and the cross of limbs leads to mesh penetration.

### 25 Mar 2024 v3

@@ 3.5 Ablation Study
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of "2 param$+$1 vert" has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
4 Discussion
Of course, W-HMR is still not a perfect model, and for this reason, we show the failure cases in Fig. 9. We can find that our model is still not able to achieve accurate reconstruction when facing extreme poses (yoga, parkour), which may be due to the lack of relevant training data or may be limited by the iterative method. Secondly, when the human limbs are close to the body, there is often through-molding, which requires an extra optimization-based method to improve it or a human model with a higher degree of freedom. Third, W-HMR seems helpless against some severely distorted images taken at extremely rare angles (e.g., the image in the third row from the left). After all, as a weakly supervised method, W-HMR is limited in its ability to predict focal length in the absence of rich annotations for training. It is also true that W-HMR is unable to correctly establish the relationship between the person and the ground for images that lack background references. These will be one of the directions worth researching in the future.
Another important direction for human mesh recovery in world space is to capture the global trajectory of 3D human bodies. W-HMR mainly solves the problems caused by camera position and rotation. Although W-HMR can recover the human body posture in world space, it cannot obtain the absolute spatial position of the human body. Absolute spatial position is essential for the practical application of motion capture. However, at present, most related research still needs to rely on additional equipment to assist. This will be an important future research direction for motion capture based on monocular images.
Figure 9: Failure cases on unusual samples. Although our model performs well in most cases, W-HMR still fails in some extreme situations. For example, humans have difficult poses, body parts are hard to distinguish from the background, and the cross of limbs leads to mesh penetration.
5 Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. "Weak-supervised" solves the problem of insufficient focal length labels. "Camera calibration" solves the problem caused by image distortion. The focal length predictor can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels, which are more suitable for human recovery tasks. Additionally, we introduced a new orientation correction module that is simpler and more robust than previous methods. Moreover, we proved that it can be ported to other models and work. With a carefully designed training paradigm and model framework, we achieved stable training and impressive performance across multiple datasets in both coordinates. We hope that our work will contribute to the study of human mesh recovery in world space. The codes and related materials will be kept updated. Everyone is welcome to acquire them from our project page https://yw0208.github.io/w-hmr/.